OpenStudy (anonymous):

Find the derivative of the given function 2+ sqrt (x) divided by x^2..... we haven't learned quotient rule yet... *answer is - (8+3sqrt(x) divided by (2x^3))

OpenStudy (anonymous):

\[2+\sqrt{x}\div x^2\]

OpenStudy (anonymous):

[2+sqrt(x)]/x^2 = (2/x^2) + [sqrt(x)/x^2] = 2x^-2 + x^(-3/2) Use simple differentiation d/dx ( 2x^-2) = -4x^-3 = -4/(x^3) d/dx (x^-3/2) = (-3/2)*[x^(-5/2)] thus adding them you get, -4/(x^3) + (-3/2)*[x^(-5/2)] = -4/(x^3) - [3/(2*sqrt(x^5)] Using Least Common Multiple, you get (-8+3sqrt(x))/(2x^3) Lemme know if you didnt understand..i'll write it on a paper.

OpenStudy (anonymous):

yes could you please write it on paper